# Yong Ji1, Junye Li1 and Rui Yang∗2,3

### v3

@@ 2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of specification property in dynamical systems.
@@ 2.1 Three types of metric mean dimensions
Let $(X,d,f)$ be a t.d.s.. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on the space $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}\{d(f^{i}x,f^{i}y)\},$
@@
$B_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ =$=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
@@
Let $Z$ be a nonempty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$. Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
@@
Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 00, i.e.,
@@
Specifically, topological entropy fails to distinguish different levels of topological complexity within the infinite entropy systems
Analogous to the definition of box dimension in fractal geometry, we formulate the refined entropy-like quantity for infinite entropy systems using $\epsilon$-upper capacity topological entropy.
@@
In the case where $Z$ is equal to $X$, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$ respectively. Then the upper capacity metric mean dimension of $X$ is reduced to Lindenstrauss-Weiss’s definition introduced in [LW00]. Such a formulation extends the concept of the metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant.
Notice that the (upper capacity) metric mean dimension is defined by detecting how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon$ tends to 00. That is, for sufficiently small $\epsilon>0$, we may think of
@@
It is straightforward to observe that any system with finite topological entropy has zero metric mean dimension. This, in turn, shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
The classical Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogue of the Hausdorff dimension and packing dimension in fractal geometry. The corresponding dimensional characterizations of metric mean dimensions, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the Bowen and packing metric mean dimensions. One may refer to [BR23, CLS21, LL24, FKO24, YCZ25] for the results involving Bowen and packing metric mean dimensions.
For $Z\subset X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
@@
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases, this ensures that the limit
@@
exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
$h_{top}^{B}(f,Z,\epsilon)=\inf\{s>0:M(Z,s,\epsilon)=0\}=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
@@
For $Z\subseteq X$, $s\in\mathbb{R}$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
@@
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So some modifications are needed to obtain the property:
@@
There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 00. That is, the critical value is given by
@@
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$. We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23b, Proposition 2.6].
@@
$(3)$ If $Z$ is a nonempty subset of $X$, then for every $10<\epsilon<$ 1, we have
@@
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$. Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions are equal.
Given a subset of phase space, it is not easy to calculate its precise Bowen and packing metric mean dimensions. However, in the multifractal analysis of dynamical systems the Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets. Consequently, the powerful tool yields the lower bounds of metric mean dimension of the fractal-like sets. $\epsilon$-Bowen and packing topological entropies of Entropy Distribution Principles are stated as follows:
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and for every $n$ and for each Bowen ball $B_{n}(x,\epsilon)$ with $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, one has $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$, then $h_{top}^{B}(f,Z,\epsilon)\geq s$.
@@
Let $(X,d,f)$ be a t.d.s. and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $s\geq 0$. If there exist a $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ that converges to $\infty$ satisfying $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then $h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s$.
@@
By definition, it is enough to prove that for sufficiently small $\delta>0$,
@@
holds for any Borel subset $E\subset Z$ with $\mu(E)>0$.
For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of the set $E$. Then, by $5r$-covering lemma stated in [Mat95, Theorem 2.1], there is a countable pairwise disjoint subcollection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, and in particular the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ Hence, one has
@@ 2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
Let $C(X)$ denote the space consisting of real-valued continuous functions defined on $X$, equipped with the supremum norm. By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
@@
where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. It follows from the weak-topology on $M(X)$ that the natural embedding from $X$ to $M(X)$, given by $x\mapsto\delta_{x}$, is continuous. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the sense of the weak-topology.
Given an invariant measure $\mu\in M_{f}(X)$, a finite family $\xi=\left\{C_{1},\cdots,C_{k}\right\}$ of Borel subsets of $X$ is a measurable partition of $X$ if $\mu(C_{i}\cap C_{j})=0$ for all $i\neq j$ and $\mu(X\setminus\cup_{i=1}^{k}C_{i})=0$. Then the partition entropy of $\xi$ w.r.t. $\mu$ is defined by
$H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$
We define the measurable partition of $\mu\in M_{f}(X)$ w.r.t. $\xi$ as
@@
where $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the $n$ partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
@@
where the supremum is taken over all finite Borel measurable partitions of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimensions is the measure-theoretic metric mean dimensions of invariant measures by measuring the divergent rates of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic treatment about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25], and the divergent rates of different types of measure-theoretic $\epsilon$-entropies coincide in terms of ergodic measures (cf. [YCZ25, Theorem 1.1]). Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
@@
    The Kolmogorov-Sinai $\epsilon$-entropy (cf. [GS21]) of $\mu\in M_{f}(X)$ is given by
    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi).$
@@
    The lower Katok $\epsilon$-entropy of $\mu$ is given by
@@
[S22, Lemma 4.1] Let $\mu\in M_{f}^{e}(X)$ and $\mathcal{U}$ be a finite open cover of $X$ with ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and ${\rm Leb}(\mathcal{U})\geq\epsilon$. Then for any $\delta\in(0,1)$, one has
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
Consequently, $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
The following variational principles verify the relations between the metric mean dimension and the two types of measure-theoretic $\epsilon$-entropies.
@@
(cf.[GS21, Theorem 3.1, Remark 3.6], [S22, Proposition 7.3]) Let $(X,f)$ be a t.d.s. with a metric $d$. Then for any $\delta\in(0,1)$,
@@
The results are also valid for lower metric mean dimension.
A nice survey involving the variational principles of metric mean dimension in terms of different types of measure-theoretic $\epsilon$-entropies [LW00, GS21, S22] is summarized in [YCZ25, Theorems 1.2 and 1.3].
2.3 Specification property
Sometimes, it is difficult to find a real orbit with desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as specification property.

### v4

@@ 2 Preliminaries
In this section, we recall the concepts of metric mean dimensions in both topological and measure-theoretic situations, and state the precise definition of the specification property in dynamical systems.
@@ 2.1 Three types of metric mean dimensions
Let $(X,d,f)$ be a TDS. For any natural number $n\in\mathbb{N}$, we define the $n$-th Bowen metric $d_{n}$ on $X$ as
$d_{n}(x,y):=\max_{0\leq i\leq n-1}d(f^{i}x,f^{i}y),$
@@
$B_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)<\epsilon\},$
$\overline{B}_{n}(x,\epsilon):$ $=\{y\in X:d_{n}(x,y)\leq\epsilon\}$
@@
Let $Z$ be a non-empty subset of $X$. Fix $\epsilon>0$. A set $E\subset Z$ is said to be an $(n,\epsilon)$-separated set of $Z$ if any distinct $x,y\in E$ implies $d_{n}(x,y)>\epsilon$. Denote by $s_{n}(Z,\epsilon)$ the largest cardinality of $(n,\epsilon)$-separated sets of $Z$. The $\epsilon$-upper capacity topological entropy of $Z$ is given by
@@
Recall that the upper capacity topological entropy of $Z$ is the limit of $\epsilon$-upper capacity topological entropy of $Z$ as $\epsilon$ goes to 0, i.e.,
@@
To get more information about the dynamics of infinite entropy systems, analogous to the definition of box dimension in fractal geometry, we formulate a refined entropy-like quantity using the $\epsilon$-upper capacity topological entropy.
@@
If $Z=X$, the upper capacity metric mean dimension of $X$ is reduced to the metric mean dimension introduced by Lindenstrauss-Weiss [LW00]; in this case, we use the notation $h_{top}(f,X,\epsilon)$ instead of $h_{top}^{UC}(f,X,\epsilon)$, and $\overline{\rm mdim}_{M}(f,X,d)$ instead of $\overline{\rm mdim}_{M}^{UC}(f,X,d)$, respectively. Such a formulation extends the concept of metric mean dimension to any subsets of the phase space, which is not necessarily compact or $f$-invariant. Notice that the (upper capacity) metric mean dimension is defined by measuring how fast the $\epsilon$-upper capacity topological entropy diverges to $\infty$ as $\epsilon\to 0$. That is, for sufficiently small $\epsilon>0$, we may think of
@@
It is easy to see that any system with finite topological entropy has zero metric mean dimension. This shows that the metric mean dimension is a suitable “measure” for capturing the dynamics of the infinite entropy systems.
The Bowen topological entropy[Bow73] and packing topological entropy[FH12] in dimension theory, which are defined by Carathéodory-Pesin structures, are the analogues of the Hausdorff dimension and packing dimension in fractal geometry. The corresponding dimensional characterizations for metric mean dimension, which have been introduced in [LV21, YCZ22, W21] through Carathéodory-Pesin structures, are the following Bowen and packing metric mean dimensions.
For $Z\subset X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
@@
where the infimum is taken over all finite or countable families $\{B_{n_{i}}(x_{i},\epsilon)\}_{i\in I}$ such that $Z\subset\cup_{i\in I}B_{n_{i}}(x_{i},\epsilon)$ with $x_{i}\in X,n_{i}\geq N$ for all $i\in I$.
The quantity $M(Z,s,N,\epsilon)$ is non-decreasing as $N$ increases. So the limit
@@
exists. There is a critical value of the parameter $s$, which we call $\epsilon$-Bowen topological entropy, such that $M(Z,s,\epsilon)$ jumps from $\infty$ to 0, i.e.,
$h_{top}^{B}(f,Z,\epsilon)$ $=\inf\{s>0:M(Z,s,\epsilon)=0\}$
$=\sup\{s>0:M(Z,s,\epsilon)=\infty\}.$
@@
For $Z\subseteq X$, $s>0$, $N\in\mathbb{N}$ and $\epsilon>0$, we put
@@
Observe that the quantity $P(Z,s,\epsilon)$ does not satisfy the countable additive property with respect to $Z$. So the modifications are needed to obtain this property:
@@
There is a critical value of the parameter $s$, which we call $\epsilon$-packing topological entropy, such that $\mathcal{P}(Z,s,\epsilon)$ jumps from $\infty$ to 0. That is, the critical value is given by
@@
One can replace $\limsup_{\epsilon\to 0}$ by $\liminf_{\epsilon\to 0}$ in above definitions to obtain the corresponding three types of lower metric mean dimensions on subsets, which are denoted by $\underline{\rm mdim}_{M}^{S}(f,Z,d)$ with $S\in\{UC,B,P\}$. If $\underline{\rm mdim}_{M}^{S}(f,Z,d)=\overline{\rm mdim}_{M}^{S}(f,Z,d)$, we call the common value ${\rm mdim}_{M}^{S}(f,Z,d)$ the (capacity/Bowen/packing) metric mean dimension of $Z$.
We summarize some fundamental properties of metric mean dimensions derived in [YCZ22, Proposition 3.4] and [YCZ23, Proposition 2.8].
@@
$(3)$ If $Z$ is a non-empty subset of $X$, then for every $0<\epsilon<1$, we have
@@
Consequently, $\overline{\rm mdim}_{M}^{B}(f,Z,d)\leq\overline{\rm mdim}_{M}^{P}(f,Z,d)\leq\overline{\rm mdim}_{M}^{UC}(f,Z,d)$.
Furthermore, if $Z$ is $f$-invariant and compact, then the three types of metric mean dimensions coincide.
Given a certain subset of phase space, in general it is not easy to calculate its precise Bowen and packing metric mean dimensions. The Entropy Distribution Principles allow us to get a lower bound for $\epsilon$-Bowen and $\epsilon$-packing topological entropies of certain fractal-like sets, and hence yields the lower bounds of metric mean dimension of the fractal-like sets.
The Entropy Distribution Principles of $\epsilon$-Bowen and packing topological entropies are stated as follows:
@@
[BR23, Lemma 13] Let $(X,d,f)$ be a TDS and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$. If there exist $\mu\in M(X)$ and a constant $C>0$ such that $\mu(Z)>0$ and $\mu(B_{n}(x,\epsilon))\leq Ce^{-ns}$ for every each Bowen ball $B_{n}(x,\epsilon)\cap Z\neq\emptyset$, then
$h_{top}^{B}(f,Z,\epsilon)\geq s.$
@@
Let $(X,d,f)$ be a TDS and $\epsilon>0$. Suppose that $Z$ is a Borel set of $X$ and $0<s<\infty$. If there exist $\mu\in M(X)$, a constant $C>0$ and a strictly increasing sequence $\{n_{i}\}\subset\mathbb{N}$ such that $\mu(Z)>0$ and $\mu(B_{n_{i}}(x,\epsilon))\leq Ce^{-n_{i}s}$ for each $i$ and for any $x\in Z$, then
$h^{P}_{top}(f,Z,\frac{\epsilon}{6})\geq s.$
@@
By definition of the $\epsilon$-packing topological entropy, it is enough to prove that for every sufficiently small $\delta>0$,
@@
for any Borel subset $E\subset Z$ with $\mu(E)>0$.
For each $i$, consider the cover $\mathcal{A}_{i}=\{B_{n_{i}}(x,\frac{\epsilon}{5}):x\in E\}$ of $E$. Then, by $5r$-covering lemma [Mat95, Theorem 2.1], there exists a countable pairwise disjoint sub-collection $\{B_{n_{i}}(x_{j},\frac{\epsilon}{5})\}_{j\in J}$, where $J$ is a index set, such that $E\subset\bigcup_{j\in J}B_{n_{i}}(x_{j},\epsilon).$ In particular, the collection of closed Bowen balls $\{\overline{B}_{n_{i}}(x_{j},\frac{\epsilon}{6})\}_{j\in J}$ is pairwise disjoint. Hence, one has
@@ 2.2 Measure-theoretic $\epsilon$-entropy of invariant measures
Let $C(X)$ denote the space of real-valued continuous functions on $X$, equipped with the supremum norm $||\cdot||$. By virtue of the Riesz representation theorem for compact metric spaces, we are able to endow $M(X)$ with the weak-topology, which is metrizable by the following metric:
@@
where $\{\phi_{k}\}$ is a countable family of dense subsets of $C(X)$ taking values in $[0,1]$. Thus, the diameter of $M(X)$ in the sense of the metric $D$ is always less than $1$, and for fixed $\mu,\nu\in M(X)$, both $D(\mu,\cdot)$ and $D(\cdot,\nu)$ are convex functions on $M(X)$. Furthermore, it is well-known that $M(X)$ and $M_{f}(X)$ are compact convex sets in the weak-topology.
Given $\mu\in M_{f}(X)$ and a finite Borel partition $\xi=\left\{C_{1},\cdots,C_{k}\right\}$, the partition entropy of $\xi$ w.r.t. $\mu$ is defined by $H_{\mu}(\xi)=-\sum_{i=1}^{k}\mu(C_{i})\log(\mu(C_{i})).$ We define the measure-theoretic entropy of $\mu$ w.r.t. $\xi$ as
@@
Denote by $\xi^{n}:=\vee_{i=0}^{n-1}f^{-i}\xi$ is the $n$-th join of the preimage partitions $\xi,f^{-1}\xi,\ldots,f^{-(n-1)}\xi$ on $X$. The measure-theoretic entropy of $\mu$ is defined as
@@
where the supremum ranges over all finite Borel partitions of $X$.
For topological dynamical systems, the measure-theoretic counterpart of metric mean dimension is the measure-theoretic metric mean dimension of invariant measures. It measures the divergent rate of the limits of the measure-theoretic $\epsilon$-entropies such as Brin-Katok $\epsilon$-entropy, Katok $\epsilon$-entropy, Kolmogorov-Sinai $\epsilon$-entropy, and the others. A systemic study about the measure-theoretic $\epsilon$-entropies is stated in [YCZ25]. Here, we only recall two types of the measure-theoretic $\epsilon$-entropies.
Let $\mu\in M_{f}(X)$ and $\epsilon>0$.
@@
    The Kolmogorov-Sinai $\epsilon$-entropy [GS21] of $\mu$ is given by
    $\inf_{{\rm diam}(\xi)<\epsilon}h_{\mu}(f,\xi),$
    where the infimum is taken over all finite Borel partitions of $X$ with diameter less than $\epsilon$.
@@
    The Katok $\epsilon$-entropy of $\mu$ [Kat80] is defined by
@@
[S22, Lemma 4.1] Let $(X,d,f)$ be a TDS and $\mu\in M_{f}^{e}(X)$, If $\mathcal{U}$ is a finite open cover of $X$ with the diameter ${\rm diam}(\mathcal{U})\leq 4\epsilon$ and the Lebesgue number ${\rm Leb}(\mathcal{U})\geq\epsilon$, then for any $\delta\in(0,1)$,
$\underline{h}_{\mu}^{K}(f,4\epsilon,\delta)\leq\inf_{\xi\succ\mathcal{U}}h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta),$
where the infimum $\xi$ ranges over all finite Borel partitions of $X$ refining than $\mathcal{U}$.
Consequently, for every $\epsilon>0$, one has $\underline{h}_{\mu}^{K}(f,16\epsilon,\delta)\leq\inf_{{\rm diam}(\xi)<4\epsilon}\limits h_{\mu}(f,\xi)\leq\underline{h}_{\mu}^{K}(f,\epsilon,\delta).$
We state the variational principles between the metric mean dimension and the above two types of measure-theoretic $\epsilon$-entropies.
@@
(cf. [YCZ25, Theorems 1.2 and 1.3]) Let $(X,d,f)$ be a TDS. Then for any $\delta\in(0,1)$,
@@
The results are also valid for $\underline{\rm mdim}_{M}(f,X,d)$ by changing $\limsup_{\epsilon\to 0}$ into $\liminf_{\epsilon\to 0}$.
2.3 Specification property of TDSs
In general, it is difficult to find a real orbit with the desired dynamical behavior, while one can get an approximate orbit of a point that $\epsilon$-shadows any specified finite orbit-segments, which is known as the specification property in TDSs.